OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (wolfe8):

I'm guessing you are confused about the absolute symbol? Well, when you see a number in it, for example |y|, it means that when you solve it it will be either -y or +y. Here, your absolute notation encloses x+7. What does that mean?

OpenStudy (anonymous):

I have no clue :P

OpenStudy (wolfe8):

Alright. Tell me how you would solve it if those were normal brackets.

OpenStudy (solomonzelman):

-3|x + 7| = -12 divide both sides by a negative 3 |x + 7|=4 So.... x+7=-4 OR x+7=4 lets work the first equation x+7=-4 x=-4-7=-13 OR (let work the:) x+7=4 x=4-7=-3 So the answers are -3 and -13! GOOD?

OpenStudy (solomonzelman):

Still don't understand? Ask if you don't....!

OpenStudy (anonymous):

I understand thanks!

OpenStudy (solomonzelman):

Medal, Please?

OpenStudy (anonymous):

Solomon, I owed you a medal! There you go debt-payment!

OpenStudy (solomonzelman):

Thanks Bro! Good Bye!